### v2

@@ IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted analysis of deceptive control. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE), and (2) additional perspectives captured by adopting formative metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of deceptive control. As suggested above, these two metrics can form three partitions explaining the different levels of capacity regarding deceptive control. Using these partitions, we can determine the best model for balancing two sub-abilities of deceptive control; GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of deceptive control was Llama. Although Llama performed better against weak citizens than strong citizens, Llama player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. So, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of deceptive control behavior.
Second, proposed metrics provide additional perspectives on investigating deceptive control, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, win rate (WR) and living round (LR). For example, both GPT-3.5 and Llama showed relatively low WR scores, but regarding LR, they achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that Llama outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at controlling information, but Llama is not. This example shows that our proposed metrics are more suitable for examining deceptive control ability than existing summative metrics.
Similarly, our proposed metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at deceptive control regarding VR and VE.
V Qualitative Analysis
@@
To reconfirm the findings of quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23]. In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
V-A Analysis Procedure
@@
Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs. To enrich our quantitative analysis results, we adopted a bottom-up inductive approach to discover the underlying patterns in the data. The detailed analysis process was as follows. First, we defined the initial labels of the abnormal patterns from 15 randomly selected gameplay logs. Three annotators with computer science backgrounds participated in this process. They read the logs to understand the spy’s gameplay situation when establishing initial labels. To avoid bias towards a specific LLM, we hid the information of the models in the process. Second, we refined these initial labels through an iterative process. Because LLMs perform one or two reasoning steps to decide on an action during the gameplay of SpyGame, a gameplay log is a series of several reasoning steps. So, we used one step as the unit for this analysis. We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated. The agreement was measured using Fleiss’ Kappa[24], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement. Finally, using these settled labels, we labeled 506 reasoning steps from 168 games. Consequently, we identified three major categories and five subcategories.
@@
TABLE IV: Frequency of each category found in qualitative analysis
$N$* Character ambiguity Memory distortion Dissociation
@@ V-B Resulting Categories
This section illustrates the results of our thematic analysis. For each category, we explain its definition, report its frequency, and provide some examples. Table IV summarizes the frequencies of all categories.
@@ V-B2 Dissociation
Llama2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[25], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes [26]. Thus, LLM players can misinterpret their words as those of others in the log, which triggers dissociation.
@@ V-B3 Memory distortion
GPT-3.5 (Spy): Based on the conversation between Player2 and Player3, it seems that the location is likely ...
@@ V-B4 Character ambiguity
Character ambiguity is a case in which the spy confuses their own identity with that of a citizen. This category is named after a similar psychological concept [28]. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of character ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with character ambiguity.
@@
Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
@@
Llama2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
@@
We suspect that this category occurs is due to the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering character ambiguity.
V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: subtext inference and deceptive control.
First, memory distortion and dissociation may hinder a spy from inferring subtext during the game. When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in subtext inference. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate. Similarly, when dissociation occurs, the spy regards apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
@@ V-C Reconfirmation of Quantitative Analysis
Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control. When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics. Also, Llama showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for deception. When character ambiguity occurs in the spy’s reasoning steps, the spy confuses their character with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
@@ VI Conclusion
To assess whether LLMs can effectively support obscured communication, we addressed two key limitations in prior evaluation methodologies. The first is the reliance on coarse-grained quantitative metrics, which fail to capture fine-grained communicative behaviors. The second is the lack of structured qualitative analyses that can contextualize evaluation results or uncover the reasoning processes underlying model behavior. To investigate these limitations in a controlled setting, we conducted experiments using SpyGame, a variant of Spyfall, and collected gameplay logs from four LLMs. We analyzed these logs through both quantitative and qualitative methods. First, we proposed eight fine-grained metrics designed to evaluate two core abilities: subtext inference and deceptive control. These metrics demonstrated more effective in identifying subtle behavioral differences between models. Second, we performed a thematic analysis to identify reasoning failures not captured by quantitative metrics. This analysis revealed four major categories of failure that align with the patterns observed in the quantitative evaluation, offering complementary insights. Our findings highlight the importance of microscopic, multi-perspective evaluation for understanding how LLMs perform in obscured communicative settings. We hope this study contributes to a deeper understanding on LLMs in complex social interactions.

### v3

@@ IV-C3 Discussion
As aforementioned, we found no model achieved a statistically significant advantage, except for Guess Success (GS) when against strong citizens. Nevertheless, several cautious insights emerge from the patterns observed in experimental results. In this section, we outline two aspects of these findings: (1) the additional perspective gained from fine-grained deception metrics versus coarse-grained metrics, and (2) differences in how LLMs manage suspicion.
First, the formative metrics provide a complementary perspective beyond outcome-level summative metrics. Specifically, evaluating based solely on the coarse-grained metrics, illustrated in V(a) and VI(a) (WR and LR), may mislead conclusions about an LLM’s ability to support a core aspect of obscured communication. For instance, an evaluation based solely on WR and LR may misleadingly suggest that Llama2 demonstrates superior performance to GPT-3.5 across experiments. In contrast, our fine-grained analysis offers a more nuanced interpretation: Llama2 recorded exceptionally high Caught Rate (CR) scores of 100% and 60% against strong and weak citizens, respectively. These findings reflect a deficiency in Llama’s capacity to support deceptive control, as the model was frequently subjected to decisive levels of suspicion, resulting in direct defeat.
A similar misinterpretation arises when comparing Gemini and Llama using coarse-grained metrics. Although both models achieved comparable WR, Llama exhibited a much higher average LR, which might misleadingly imply that the model succeeded in deceiving citizens and thereby extended its survival. However, our metrics show that Llama also had a substantially higher CR, meaning it was frequently the primary target of suspicion. This implies that more prolonged survival may not necessarily reflect successful deception, but rather passive or non-committal behavior that delays elimination without effectively evading suspicion. Such cases highlight a potential risk in using metrics like living round as reward signals in RL-based optimization, as they could incentivize specification gaming by encouraging models to optimize for the metric without actually demonstrating the intended deceptive behavior[25].
Second, the multifaceted interpretation framework presented in Table VII enables a multifaceted interpretation of how models manage suspicion during gameplay. Although the differences are not statistically significant, each model exhibits a distinct pattern across these two dimensions. GPT-4 and GPT-3.5 demonstrate relatively low VR and high VE, aligning with the ideal pattern where the spy both avoids being directly targeted and induces dispersed suspicion. In contrast, Llama shows a higher VR and lower VE, indicating concentrated suspicion and a lack of effective misdirection. Gemini is assigned to the moderate case, maintaining a low VR but struggling to distribute suspicion broadly. These results suggest that while coarse metrics obscure such differences, formative metrics can reveal the specific ways in which each model either succeeds or struggles to support deceptive control.
V Qualitative Analysis via Thematic Coding
@@
To investigate RQ2 regarding the reasoning errors that hinder LLMs, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning when deciding their actions by coding abnormal patterns in 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[26]. Furthermore, these reasoning errors appear to explain concrete failure modes observed in our quantitative metrics, particularly in Caught Rate (CR) and Guess Success Rate (GSR). This alignment between qualitative and quantitative findings helps answer RQ2 by showing not only what types of reasoning errors occur, but also how they directly undermine the LLM’s ability to avoid detection or correctly infer hidden information.
V-A Qualitative Coding Procedure
@@
Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs. To enrich our quantitative analysis results, we adopted a bottom-up inductive approach to discover the underlying patterns in the data. The detailed analysis process was as follows. First, we defined the initial labels of the abnormal patterns from 15 randomly selected gameplay logs. Three annotators with computer science backgrounds participated in this process. They read the logs to understand the spy’s gameplay situation when establishing initial labels. To avoid bias towards a specific LLM, we hid the information about the models in the process. Second, we refined these initial labels through an iterative process. Because LLMs perform one or two reasoning steps to decide on an action during the gameplay of SpyGame, a gameplay log is a series of several reasoning steps. So, we used one step as the unit for this analysis. We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated. The agreement was measured using Fleiss’ Kappa[27], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement[28]. Finally, using these settled labels, we labeled 506 reasoning steps from 168 games. Consequently, we identified three major categories and five subcategories.
@@
TABLE VIII: Frequency of each category found in qualitative analysis
NN* Character ambiguity Memory distortion Dissociation
@@ V-B Resulting Categories
This section illustrates the results of our thematic analysis. For each category, we explain its definition, report its frequency, and provide some examples. Table VIII summarizes the frequencies of all categories.
@@ V-B2 Dissociation
Llama2 (Spy): ‘Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...’
We suspect that this category occurs because LLMs struggle to differentiate between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[29], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes [30]. Thus, LLM players can misinterpret their words as those of others in the log, which triggers dissociation.
@@ V-B3 Memory distortion
GPT-3.5 (Spy): ‘Based on the conversation between Player2 and Player3, it seems that the location is likely ...’
@@ V-B4 Character ambiguity
Character ambiguity is a case in which the spy confuses their own identity with that of a citizen. This category is named after a similar psychological concept [32]. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of character ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the following paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with character ambiguity.
@@
Gemini (Spy): ‘Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...’
@@
Llama2 (Spy): ‘... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...’
@@
We suspect that this category occurs due to the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules holds more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters described in the game rules, triggering character ambiguity.
V-C Discussion
In this section, we first discuss how the reasoning error categories identified via thematic analysis help interpret the performance differences in subtext inference and deceptive control observed in our quantitative metrics. Then, we extend this discussion by exploring the potential generalizability of these error categories to broader LLM applications beyond the game-based setting.
TABLE IX: Linking reasoning failures to quantitative performance
Core Aspect Reasoning Error Behavioral or Reasoning Effect Quantitative Effect
Subtext Inference Memory Distortion & Dissociation Incorrect deduction of hidden location, leading to failure of Guessing action lower GSR
lower ICR
lower IDR
Deceptive Control Character Ambiguity Increasing suspicion from other players, leading to failure of misleading opponents higher CR
higher VR
lower VE
As illustrated in Table IX, we map the identified reasoning errors to their quantitative impacts on the reasoning processes or behaviors of LLMs during the game. First, memory distortion and dissociation may hinder a spy from inferring subtext during the game. When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the frequency trend was similar to that of the metrics. For example, against strong citizens, GPT-3.5 exhibited the highest number of memory distortion and dissociation errors, and also showed the lowest performance across the subtext inference metrics. Similarly, GPT-4 exhibited fewer memory distortion and dissociation errors, along with higher subtext inference performance compared to the other models. We interpret this alignment as stemming from the quality of information the model considers during reasoning. To infer the hidden location, the spy should capture or synthesize clues from the conversation. However, when memory distortion occurs, the spy may misremember or misinterpret prior utterances as valid sources. Likewise, in the case of dissociation, the model may mistakenly treat its own prior statements as valid information provided by others. These misunderstandings can lead to incorrect conclusions, hindering the spy’s successful inference of the subtext.
@@ V-C Discussion
Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control. When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the frequency trend was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics. Also, Llama showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for deception. When character ambiguity occurs in the spy’s reasoning steps, the spy confuses their character with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
Further, these findings not only highlight key failure modes in obscured communication but also point to a broader applicability of these reasoning errors across diverse LLM use cases. While our experimental setup is based on a social deduction game, many components of our qualitative analysis are not inherently limited to game-based settings. For example, memory distortion can affect retrieval-augmented generation (RAG) or multi-turn QA, in which LLMs fail to retain retrieved or prior context appropriately. Likewise, character ambiguity is also relevant in role-specific instruction following, such as persona-based assistants or simulation environments, where confusion about LLMs’ intended role can result in inconsistent or constraint-violating responses. These observations suggest that the taxonomy of reasoning failures developed in this study may offer a useful framework for diagnosing and improving LLM behavior beyond the context of obscured communication.
@@ VI Conclusion
In this study, we formulated two research questions to guide our investigation into how LLMs support obscured communication. The first concern is whether prior coarse-grained outcome metrics are sufficient for evaluating the fine-grained skills required in such contexts. The second investigates the underlying reasoning failures that can be systematically identified, and how these failures hinder LLMs’ ability to perform effectively in adversarial communicative scenarios.
To examine the first question, we proposed six fine-grained quantitative metrics targeting two core dimensions of obscured communication: subtext inference and deceptive control. Through empirical evaluation using gameplay data collected from four LLMs in the SpyGame environment, we demonstrated that these metrics facilitate the diagnosis of intermediate behavioral patterns and expose the limitations of conventional evaluation methods that rely solely on outcome-level indicators such as win rate.
To address the second question, we conducted a thematic qualitative analysis of gameplay logs capturing the LLMs’ reasoning processes and actions. This analysis revealed recurring reasoning error categories, such as memory distortion, dissociation, and character ambiguity, which not only reflect systematic failure modes but also align with patterns observed in the quantitative results. This alignment provides explanatory insight into the underlying causes of LLM performance breakdowns in obscured communication tasks.
Despite our contribution, we aim to broaden its scope in future work by improving two points. First, we plan to extend and validate our analysis beyond the current game-based environment, exploring its generalizability to broader domains. In parallel, we will expand the evaluated roles beyond the single-agent spy to include diverse multi-agent settings, enabling the assessment of collaborative deception and coordination skills that are difficult to capture in the current setup.
Overall, this study contributes a structured evaluation framework that combines fine-grained metrics with qualitative error analysis, offering a deeper understanding of LLM behavior in complex communicative settings. We hope this study contributes to a deeper understanding of LLMs in complex social interactions.